OpenStudy (anonymous):

What is the domain of f(x)=3x -²

OpenStudy (solomonzelman):

\(\displaystyle 3x^{-2}\Longrightarrow \frac{3}{x^2}\) you know that right?

OpenStudy (solomonzelman):

I am using the rule: \(\LARGE\displaystyle a^{-b}=\frac{1}{a^b}\) and when I have a constant in front, such that: \(\LARGE\displaystyle c(a)^{-b}=c\cdot \frac{1}{a^b}=\frac{c}{a^b}\)